OpenStudy (love_to_love_you):

PLEASE HELP. WILL FAN, MEDAL, AND LEAVE TESTIMONY A doctor's office schedules 15-minute appointments and half-hour appointments for weekdays. The doctor limits these appointments to, at most, 30 hours per week. Write an inequality to represent the number of 15-minute appointments, x, and the number of half-hour appointments, y, the doctor may have in a week. Show all work.

OpenStudy (love_to_love_you):

@Hero @nickhunter172

OpenStudy (anonymous):

30*60=1800 (to set it equal to minutes.) Any combination can equal 1800 minutes, so to find an inequality, you have to set the number of fifteen minute and 30 minute ones less than or equal to 1800 minutes. Let x equal # of 15 minute appointments. Let y equal the # of 30 minute appointments. Therefore \[15x+30y \le 1800\] Hopefully that's right.

OpenStudy (love_to_love_you):

Yeah that looks correct

OpenStudy (anonymous):

Awesome. If you want to use hours, you can also use \[0.25x +0.5y \le 30\]

OpenStudy (love_to_love_you):

Alrighty, wanna help me with my assignment? ;)

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (love_to_love_you):

Thanks again for all of your help, I really do appreciate it.

OpenStudy (anonymous):

Yeah, sure. ;)

OpenStudy (love_to_love_you):

You are the most amazing person alive!

OpenStudy (love_to_love_you):

Activity 2: Graphing A juice bottling machine fi lls 1200 bottles per minute. • Graph the number of bottles fi lled versus the time (in minutes) the machine runs. • Is this model a direct variation? If so, what is the constant of variation?

OpenStudy (love_to_love_you):

Ignore the question marks, all of it is there.

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (anonymous):

Well, it is a direct variation, and the CoV (slope) is 1200. If you want to graph, then just graph y=1200x. Or, if you want seconds, y=60x.

OpenStudy (love_to_love_you):

How do I graph that though?

OpenStudy (love_to_love_you):

@satellite73 @amistre64

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (anonymous):

Well, do you know how to graph?

OpenStudy (anonymous):

y=1200x is very nearly a straight line.

OpenStudy (love_to_love_you):

Which one do you recommend graphing to get a better grade?

OpenStudy (anonymous):

The problem says minutes, so I'd go with y=1200x, probably.

OpenStudy (love_to_love_you):

Like that?

OpenStudy (anonymous):

Yes, if you're going with minutes.

OpenStudy (love_to_love_you):

Okay :)

OpenStudy (love_to_love_you):

Activity 3: Analyzing Suppose you manage a juice factory. One of your machines fi lls 850 bottles per minute. Each day it takes 30 min to get the machine started after an hour for maintenance. • Graph the number of bottles fi lled as a function of time that the machine runs. • Find the domain, range, and y-intercept of the graph. Interpret these numbers. • Write an equation for your graph.

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (anonymous):

i guess this assumes it runs 24/7 except for maintenance and warmup?

OpenStudy (love_to_love_you):

I guess so

OpenStudy (love_to_love_you):

@nickhunter172

OpenStudy (love_to_love_you):

Can you look at this by any chance and solve any of these problems???

OpenStudy (love_to_love_you):

whoops wrong file, you can check that if you'd like, thats my work for the problem I was asking you

OpenStudy (love_to_love_you):

Here's the right one.